Engineers want to design seats in commercial aircraft so that they are wide enough to fit 95​% of all males.​ (Accommodating 100% of males would require very wide seats that would be much too​ expensive.) Men have hip breadths that are normally distributed with a mean of 14.7 in. and a standard deviation of 1.1 in. Find Upper P 95.

That​ is, find the hip breadth for men that separates the smallest 95​% from the largest 5​%.
Answer #1

solution

Using standard normal table,

P(Z < z) = 95%

= P(Z < z) = 0.95

= P(Z < 1.64) = 0.95

z =1.64 Using standard normal z table,

Using z-score formula  

x= z * \sigma + \mu

x= 1.64*1.1+14.7

x= 16.504

x=17